### v1

@@ 4.1 Performance on the Industry Standard Benchmark TabArena
TabArena [40] is the most curated tabular benchmark, based on the largest number of candidate datasets considered, and created by open-source contributors from a wide range of institutions. It will appear at the NeurIPS 2025 Datasets & Benchmarks track and is thus most up-to-date. We follow the paper’s recommendation to benchmark on “TabArena-Lite”, which is a cheaper but representative version of the full benchmark using only one test fold. The benchmark contains a set of 51 datasets selected from 1053 to be representative of real-world tabular data. See Erickson et al. [40] for the list of datasets.
@@
Comparing the default performance of TabPFN-2.5 and TabPFNv2, we see a big leap in performance in Figure 3. In addition, looking at performance on each dataset in TabArena (TabPFNv2 compatible subset) in Figure 2, we see that TabPFN-2.5 clearly outperforms TabPFNv2 on almost all datasets, and is never much worse. In Appendix G, we detail the results on TabArena-Lite, showing the pairwise win rates of the different models, and comparing TabPFN-2.5 to other foundation models like TabICL [95] or LimiX [134].
@@ 4.2 Performance on Internal Benchmarks
Figure 7: TabPFN-as-MLP still outperforms tree-based models while having much faster inference speed than TabPFN. For baseline, light blue represents performance when tuned for 1 hour, and darker blue default performance. For TabPFN, we report default performance.
@@
4.4 Fast Inference with TabPFN-2.5-as-MLP
We benchmark TabPFN-2.5-as-MLP against tuned LightGBM, XGBoost, and CatBoost models, as well as the standard TabPFN-2.5 model, on our curated collection of internal open source datasets with less than 10k data points. Figure 7 illustrates representative test-split performance. Empirically, TabPFN-2.5-as-MLP offers competitive accuracy while reducing inference cost, making it attractive for high-throughput or resource-constrained deployment scenarios.
@@ 4.5 TabPFN for Causal Inference
  • For small datasets, enable the fitted-model cache: This is an experimental feature that trains and stores the model during .fit(), making subsequent .predict() calls fast by using a KV-Cache. It is enabled by setting the fit_mode parameter of TabPFNClassifier and TabPFNRegressor to fit_with_cache. However, with this setting classification models will consume approximately 6.1 KB of GPU memory and 48.8 KB of CPU memory per cell in the training dataset (regression models about 2525% less), thus it is currently only suitable for small training datasets. For larger datasets and CPU-based inference, we recommend the TabPFN-as-MLP/Tree output engine.

### v2

@@ 4.1 Performance on the Industry Standard Benchmark TabArena
TabArena [41] is the most curated tabular benchmark, based on the largest number of candidate datasets considered, and created by open-source contributors from a wide range of institutions. It will appear at the NeurIPS 2025 Datasets & Benchmarks track and is thus most up-to-date. In particular, it compares a large class of recent models, including tree-based models like CatBoost [97], LightGBM [61] or XGBoost [26], as well as newer deep-learning models like RealMLP [54], TabM [47], ModernNCA [134] or xRFM [12], and other Tabular Foundation Models like TabICL [98], TabDPT [78], LimiX [138], Mitra [139] or TabPFNv2 [53]. We follow the paper’s recommendation to benchmark on “TabArena-Lite”, which is a cheaper but representative version of the full benchmark using only one test fold. The benchmark contains a set of 51 datasets selected from 1053 to be representative of real-world tabular data. See Erickson et al. [41] for the list of datasets.
@@
Comparing the default performance of TabPFN-2.5 and TabPFNv2, we see a big leap in performance in Figure 3. In addition, looking at performance on each dataset in TabArena (TabPFNv2 compatible subset) in Figure 2, we see that TabPFN-2.5 clearly outperforms TabPFNv2 on almost all datasets, and is never much worse. In Appendix G, we detail the results on TabArena-Lite, showing the pairwise win rates of the different models, and comparing TabPFN-2.5 to other foundation models like TabICL [98], TabDPT [78] or LimiX [138], each time restricting ourselves to the subset of datasets compatible with these models.
@@ 4.2 Performance on Internal Benchmarks
Figure 7: TabPFN-as-MLP and TabPFN-as-TreeEns still outperform tree-based models while having much faster inference speed than TabPFN. For baseline, light blue represents performance when tuned for 1 hour, and darker blue default performance. For TabPFN, we report default performance. All scores are on from our internal benchmark, up to 10K data points.
@@
4.4 Fast Inference with our distillation engine
We benchmark our distillation engine which, given a training data set, outputs a multi-layer perceptron (TabPFN-2.5-as-MLP) or tree ensemble classifier (TabPFN-2.5-as-TreeEns), against tuned LightGBM, XGBoost, and CatBoost models, as well as the standard TabPFN-2.5 model, on our curated collection of internal open source datasets with less than 10k data points. Figure 7 illustrates representative test-split performance. Empirically, TabPFN-2.5-as-MLP and TabPFN-2.5-as-TreeEns offer competitive accuracy while reducing inference cost, making it attractive for high-throughput or resource-constrained deployment scenarios.
@@ 4.5 TabPFN for Causal Inference
  • For small datasets, enable the fitted-model cache: This is an experimental feature that trains and stores the model during .fit(), making subsequent .predict() calls fast by using a KV-Cache. It is enabled by setting the fit_mode parameter of TabPFNClassifier and TabPFNRegressor to fit_with_cache. However, with this setting classification models will consume approximately 6.1 KB of GPU memory and 48.8 KB of CPU memory per cell in the training dataset (regression models about 2525% less), thus it is currently only suitable for small training datasets. For larger datasets and CPU-based inference, we recommend the TabPFN-as-MLP/TreeEns.
@@ Appendix A Contributors
Special thanks to Samuel Müller for helpful discussions and support.